Alex earns a wage of 15 an hour if he works up to forty hours per week. however,for every hours, alex works in excess of 40 he earns a wage of 30 an hour。

Alex also faces a 20% tax rate and pays child-care expenses of 4 for every hour he works.

He also receives a child support payment of 80 per week. He has a total of 110 hours per week that he can dedicate to work and leisure.

Graph alex potential income constraint line
Answer #1

ANSWER

If Alex works for 40 hours,

Income after tax= 15*40*(1-0.2)=$480

Expense in child care= 4*40=$160

Child care support Income= $80

Total Income= 480 - 160 + 80 = $400

If Alex decided to work overtime,

Income after tax= (15*40+30*70)*(0.8)= $2160

Child care expenses= 4*110= $440

Child care support Income= $80

Total Income= 2160 - 440 + 80 =$1800

So Alex's income constarint line will look like